# MatDecompSDF: High-Fidelity 3D Shape and PBR Material Decomposition from Multi-View Images

### v1

@@ 3.1. Framework Overview
As illustrated in Figure 1, our framework takes a set of multi-view images of an object, along with their corresponding camera poses and foreground masks, as input. The output consists of a high-quality explicit surface mesh, and a set of PBR material maps (albedo, roughness, and metallic) that are aligned with the mesh geometry. Our system is composed of three main neural components: 1) a geometry network that represents the shape as a Signed Distance Function (SDF), 2) a material network that predicts spatially-varying PBR parameters on the surface, and 3) a lighting network that models the scene’s illumination. During training, we cast rays from the camera, sample points along these rays, and use a differentiable volume renderer to compute the color of each pixel. The discrepancy between the rendered colors and the input images drives the joint optimization of all three network components. This end-to-end optimization is analogous to complex inference processes seen in other domains, such as iterative context-aware graph inference for visual dialog (Guo et al., 2020).
@@ 3.2. Neural Scene Representation
We represent the scene using three distinct MLPs to effectively disentangle geometry, material, and lighting, a strategy proven more robust than using a single monolithic network (Zhang et al., 2021b).
@@ 3.2.1. Geometry as a Neural SDF
We represent the object’s geometry using a neural network $f_{geo}$ that approximates a Signed Distance Function (SDF). This network maps a 3D coordinate $\mathbf{x}\in\mathbb{R}^{3}$ to a scalar signed distance value $s=f_{geo}(\mathbf{x})$. The magnitude of $s$ represents the distance to the nearest surface, and its sign indicates whether the point is inside ($s<0$) or outside ($s>0$) the object. The object surface $\mathcal{S}$ is thus implicitly defined as the zero-level set of the SDF: $\mathcal{S}=\{\mathbf{x}\in\mathbb{R}^{3}\mid f_{geo}(\mathbf{x})=0\}$. This representation, popularized by works like NeuS (Wang et al., 2021), allows for the extraction of high-fidelity, watertight meshes, which is a significant advantage over direct volumetric representations like NeRF (Mildenhall et al., 2020). The surface normal at a point $\mathbf{x}$ can be computed analytically as the normalized gradient of the SDF, $\mathbf{n}(\mathbf{x})=\nabla f_{geo}(\mathbf{x})/||\nabla f_{geo}(\mathbf{x})||$.
@@ 3.2.2. Spatially-Varying PBR Materials
To represent the object’s material properties, we employ a second MLP, the material network $f_{mat}$. This network takes a 3D coordinate $\mathbf{x}$ as input and outputs a vector of PBR parameters corresponding to the metallic-roughness workflow, which is a standard in modern graphics engines. Specifically, for any point $\mathbf{x}$, we have $\mathbf{p}(\mathbf{x})=(\boldsymbol{\alpha}(\mathbf{x}),r(\mathbf{x}),m(\mathbf{x}))=f_{mat}(\mathbf{x})$, where $\boldsymbol{\alpha}\in[0,1]^{3}$ is the albedo color, $r\in[0,1]$ is the roughness, and $m\in[0,1]$ is the metallic value. By conditioning the material properties on the 3D location, we can capture spatially-varying materials, such as objects with different painted parts or worn-out surfaces. This explicit decomposition is crucial for downstream applications like material editing, a feature lacking in methods that only model view-dependent appearance (Poole et al., 2023; Lin et al., 2023).
@@ 3.2.3. Environment Lighting
We model the scene’s illumination as a distant environment map. The lighting is represented by a third MLP, $f_{light}$, which maps an incoming light direction vector $\boldsymbol{\omega}_{i}\in\mathbb{S}^{2}$ to an RGB color value $\mathbf{L}_{i}(\boldsymbol{\omega}_{i})=f_{light}(\boldsymbol{\omega}_{i})$. This formulation is flexible enough to represent complex, high-frequency lighting environments, which is a notable improvement over simpler models like spherical harmonics used in some earlier works (Zhou et al., 2021). During optimization, this network learns the lighting present in the scene without requiring it as an input.
@@ 3.3. Differentiable Rendering for Decomposition
The core of our method is a differentiable rendering process that connects our 3D scene representation to the 2D input images. We adopt the volume rendering principle from NeuS (Wang et al., 2021) to robustly find the surface intersection and then apply a physically-based rendering equation at that surface point.
@@ 3.3.1. Ray Marching and Surface Finding
[…]
Instead of a generic density $\sigma$, we use a weighting function $w(t)$ that is guaranteed to be a single, unbiased peak at the surface location. Following NeuS (Wang et al., 2021), we define a density-like function $\sigma(t)$ from the SDF value and then derive the weight $w(t)=T(t)\sigma(t)$. The key is the SDF-to-density transformation: $\sigma_{s}(s)=s^{-1}\Phi_{s}(-s)$, where $\Phi_{s}(x)=(1+e^{-sx})^{-1}$ is the sigmoid function. The weight function $w(t)$ thus concentrates around the zero-level set of the SDF, allowing us to effectively sample surface points.
@@ 3.3.2. Physically-Based Shading
[…]
@@
The terms $D,F,G$ are standard functions of the surface roughness $r$, view/light directions, and the Fresnel term $F$ additionally depends on the metallic property $m$ and albedo $\boldsymbol{\alpha}$ (Munkberg et al., 2022). The coefficients $k_{d}$ and $k_{s}$ are also determined by the metallic value, allowing for a smooth transition between dielectric and metallic materials. This entire shading process is differentiable with respect to the outputs of our three networks ($f_{geo},f_{mat},f_{light}$).
@@ 3.4. Optimization and Loss Functions
The networks are trained end-to-end by minimizing a composite loss function. The ability to learn from complex, multi-modal, and sometimes noisy data is crucial, a challenge shared by fields as diverse as sign language production (Tang et al., 2022) and WiFi-based sensing (Wang et al., 2024).
@@ 3.4.1. Photometric Reconstruction Loss
The primary supervision signal is the photometric reconstruction loss, which measures the difference between the rendered pixel colors $\mathbf{C}(\mathbf{r})$ and the ground-truth colors from the input images $\mathbf{C}_{gt}(\mathbf{r})$. We use a combination of an L1 loss and the LPIPS perceptual loss (Srinivasan et al., 2023):
(4) $\mathcal{L}_{photo}=\lambda_{L1}||\mathbf{C}(\mathbf{r})-\mathbf{C}_{gt}(\mathbf{r})||_{1}+\lambda_{LPIPS}\mathcal{L}_{LPIPS}(\mathbf{C}(\mathbf{r}),\mathbf{C}_{gt}(\mathbf{r})).$
@@ 3.4.2. Geometric Regularization
(5) $\mathcal{L}_{eikonal}=\mathbb{E}_{\mathbf{x}}\left(||\nabla f_{geo}(\mathbf{x})||_{2}-1\right)^{2}.$
@@
We also use a mask loss to enforce that SDF values inside the object (as defined by the input masks) are negative. This robust supervision is critical, much like how robust authentication is needed to prevent attacks in other sensing domains (Huang et al., 2024b).
@@ 3.4.3. Material and Lighting Priors
(6) $\mathcal{L}_{mat\_smooth}=\mathbb{E}_{\mathbf{x}_{s}}\left(||\nabla\boldsymbol{\alpha}(\mathbf{x}_{s})||_{1}+||\nabla r(\mathbf{x}_{s})||_{1}\right).$
@@
Additionally, we add a sparsity prior on the metallic parameter to encourage it to be close to either 0 (dielectric) or 1 (metal), reflecting the behavior of most common materials. We also regularize the total intensity and smoothness of the predicted environment light map from $f_{light}$ to prevent overly complex lighting from being ”baked” into the material albedo. The design of these priors is crucial for stable training, a principle that holds true when designing complex hierarchical systems, whether for federated learning (Liu et al., 2024b) or visual understanding (Guo et al., 2019b).
@@ 3.5. Implementation Details
Our geometry and material networks are implemented as MLPs with 8 layers, 256 hidden units, and sinusoidal positional encodings (Mildenhall et al., 2020) for the input coordinates to capture high-frequency details. The lighting network is a smaller 4-layer MLP. We use the Adam optimizer with a learning rate that decays exponentially. The entire framework is implemented in PyTorch. The training process, summarized in Algorithm 1, typically converges after 200k iterations on a single NVIDIA A100 GPU. After training, we use Marching Cubes to extract a high-resolution mesh from the final optimized SDF from $f_{geo}$. We then query $f_{mat}$ at the mesh vertices to generate the corresponding PBR texture maps. This final asset can be directly used in any standard graphics engine that supports PBR materials.
@@ 4.1.1. Datasets
We evaluate our method on both synthetic and real-world datasets to demonstrate its effectiveness and generalization capabilities. The diversity of data is crucial for robust evaluation, a principle that is vital across various domains, from training models for sign language (Guo et al., 2019b) to ensuring robust performance in WiFi-based activity recognition under interference (Huang et al., 2023a).
@@
For quantitative evaluation of both geometry and material decomposition, we use a synthetic dataset inspired by the one used in NVDIFFREC (Munkberg et al., 2022). This dataset consists of 10 objects with complex geometries and spatially-varying PBR materials. Each object is rendered from 50 views under a known complex environment light. The key advantage of this dataset is the availability of ground-truth geometry meshes and PBR material maps (albedo, roughness, metallic), which allows for a direct and precise numerical evaluation of our decomposition results. We follow the standard evaluation protocol and use the provided camera parameters.
@@
To validate our method on real-world data, we use the widely-adopted DTU dataset (jensen2014large). The DTU dataset contains multi-view images of various objects captured under controlled laboratory conditions with a fixed camera rig and several different lighting conditions. This dataset is a standard benchmark for multi-view surface reconstruction (Wang et al., 2021; Yariv et al., 2021). Since ground-truth material properties are not available for these real objects, evaluation is primarily focused on the quality of the geometric reconstruction and the visual fidelity of novel view synthesis and relighting results. The challenge of working with real-world data, where ground truth is scarce, mirrors challenges in other sensing domains that rely on unsupervised or weakly-supervised methods, such as gesture recognition from commodity signals (Gu et al., 2023a).
@@ 4.1.2. Evaluation Metrics
We adopt a comprehensive set of metrics to evaluate different aspects of our framework’s performance. The choice of metrics is critical for a fair comparison, reflecting the multi-faceted nature of the inverse rendering problem, much like how multi-objective optimization is key for efficient model compression (Fan et al., 2024).
@@
  • Novel View Synthesis: For both datasets, we evaluate the quality of rendered novel views against ground-truth images using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and the Learned Perceptual Image Patch Similarity (LPIPS) (Srinivasan et al., 2023). PSNR and SSIM measure pixel-level fidelity, while LPIPS better correlates with human perceptual judgment, a crucial aspect for tasks aiming at human-centric understanding, such as visual dialog (Guo et al., 2021) or emotion recognition (Gu et al., 2023b).
@@ 4.1.3. Baselines and Implementation Details
  • NeuS (Wang et al., 2021): A state-of-the-art method for high-quality neural surface reconstruction. This baseline represents methods that focus solely on geometry.
  • NVDIFFREC (Munkberg et al., 2022): A leading method for inverse rendering that directly optimizes a mesh representation. It is a strong baseline for joint geometry and material optimization.
  • Mat-NeuS (Zhang et al., 2022a): An extension of NeuS that incorporates material decomposition, making it a very relevant baseline for our approach.
  • IRON (Zhang et al., 2022b): Another powerful inverse rendering framework that demonstrates high-quality object reconstruction and material estimation.
@@
Our framework is implemented in PyTorch. All experiments are conducted on NVIDIA A100 GPUs with 40GB of memory. We use the Adam optimizer with an initial learning rate of $5\times 10^{-4}$, which decays exponentially over the course of training. Each model is trained for 300k iterations with a batch size of 512 rays. The loss weights are set empirically, with $\lambda_{L1}=1.0$, $\lambda_{LPIPS}=0.2$, $\lambda_{eikonal}=0.1$, and $\lambda_{mat}=0.01$. The design of an efficient and stable training pipeline is paramount, a concern shared by those developing complex distributed systems for tasks like federated learning (Liu et al., 2024b, a) or dealing with the security implications of advanced sensing technologies (Huang et al., 2024a). Our code and pre-trained models will be made publicly available to facilitate future research.
@@ 4.2. Quantitative Comparison
Table 1 provides a detailed breakdown of performance on the synthetic dataset, where ground-truth geometry and materials are available for precise evaluation. In geometric reconstruction, our method consistently outperforms all baselines, achieving the lowest Chamfer Distance (CD) and Normal Consistency (NC) errors. Compared to NeuS (Wang et al., 2021), which focuses solely on geometry, our joint optimization of shape and material yields a more accurate surface. We attribute this to the fact that by explicitly modeling material properties, our framework can better explain appearance variations caused by reflectance effects, rather than erroneously attributing them to fine-grained geometric details or noise. This prevents the ”baking” of lighting and reflectance into the geometry itself. Our geometric fidelity even surpasses that of leading inverse rendering methods like NVDIFFREC (Munkberg et al., 2022) and Mat-NeuS (Zhang et al., 2022a) which we believe is due to our robust lighting model and the comprehensive set of physical and geometric priors that regularize the optimization process. The challenge of disentangling correlated factors is a common theme in machine perception, where, for instance, separating an event signal from noise is key in audio-visual analysis (Zhou et al., 2022).
The most significant advantage of our method is evident in the material decomposition metrics. As shown in Table 1, MatDecompSDF achieves a substantially higher PSNR across all PBR maps (albedo, roughness, and metallic) compared to all baselines. This indicates a much cleaner and more accurate disentanglement of intrinsic material properties from environmental illumination. While methods like IRON (Zhang et al., 2022b) can recover plausible materials, they often exhibit minor artifacts, such as residual lighting baked into the albedo map. Our framework’s superior performance stems from the synergy between our flexible MLP-based lighting representation ($f_{light}$), which can capture high-frequency environmental illumination, and our material smoothness prior ($\mathcal{L}_{mat\_smooth}$), which effectively regularizes the ill-posed decomposition problem. This demonstrates that a well-constrained optimization, guided by physical principles, is paramount, a lesson that holds true in fields like model compression where multi-objective approaches are needed to balance competing goals (Fan et al., 2024).
@@
Table 1. Quantitative comparison on the synthetic dataset. We report Chamfer Distance (CD, mm, $\downarrow$), Normal Consistency (NC, degrees, $\downarrow$), and PSNR (dB, $\uparrow$) for material maps and rendering quality. Best results are in bold.
@@
NeuS (Wang et al., 2021) 0.85 16.2 - - - 29.81 -
NVDIFFREC (Munkberg et al., 2022) 0.79 15.8 25.12 23.45 24.88 31.23 28.95
Mat-NeuS (Zhang et al., 2022a) 0.75 15.1 26.03 24.11 25.01 31.55 29.24
IRON (Zhang et al., 2022b) 0.81 15.5 25.88 23.97 24.92 31.40 29.08
@@
Table 3 shows the performance on the challenging real-world DTU dataset. Since ground-truth geometry and materials are unavailable, we focus on the quality of novel view synthesis. Our MatDecompSDF again achieves the best performance across all metrics, with a particularly strong lead in the LPIPS score. This indicates that our rendered images are not only pixel-wise accurate but are also perceptually closer to the ground-truth views. We attribute this to two factors: first, the inclusion of the $\mathcal{L}_{LPIPS}$ term in our photometric loss guides the optimization towards perceptually plausible solutions, a strategy whose importance is recognized in generative modeling (Srinivasan et al., 2023). Second, by modeling the scene with physically accurate materials and lighting, our framework produces more realistic shading and highlights, which are key components of perceptual quality. This is in contrast to geometry-only methods like NeuS, whose view-dependent appearance model can sometimes produce results that look flat or lack realism under complex lighting. The ability to perform robustly on real-world data with its inherent noise and complexities is a testament to our framework’s design, echoing the need for robust systems in other real-world sensing applications, from federated learning in heterogeneous networks (Liu et al., 2024a) to secure WiFi-based gesture recognition (Gu et al., 2023a).
@@
NeuS (Wang et al., 2021) 25.31 0.871 0.145
NVDIFFREC (Munkberg et al., 2022) 25.88 0.885 0.131
Mat-NeuS (Zhang et al., 2022a) 26.05 0.892 0.124
IRON (Zhang et al., 2022b) 25.97 0.888 0.129

### v2

@@ 3.1. Framework Overview
As illustrated in Figure 1, our framework takes a set of multi-view images of an object, along with their corresponding camera poses and foreground masks, as input. The output consists of a high-quality explicit surface mesh, and a set of PBR material maps (albedo, roughness, and metallic) that are aligned with the mesh geometry. Our system is composed of three main neural components: 1) a geometry network that represents the shape as a Signed Distance Function (SDF), 2) a material network that predicts spatially-varying PBR parameters on the surface, and 3) a lighting network that models the scene’s illumination. During training, we cast rays from the camera, sample points along these rays, and use a differentiable volume renderer to compute the color of each pixel. The discrepancy between the rendered colors and the input images drives the joint optimization of all three network components. This end-to-end optimization is analogous to complex inference processes seen in other domains, such as iterative context-aware graph inference for visual dialog .
@@ 3.2. Neural Scene Representation
We represent the scene using three distinct MLPs to effectively disentangle geometry, material, and lighting, a strategy proven more robust than using a single monolithic network (zhang2021physg).
@@ 3.2.1. Geometry as a Neural SDF
[…]
@@ 3.2.2. Spatially-Varying PBR Materials
To represent the object’s material properties, we employ a second MLP, the material network $f_{mat}$. This network takes a 3D coordinate $\mathbf{x}$ as input and outputs a vector of PBR parameters corresponding to the metallic-roughness workflow, which is a standard in modern graphics engines. Specifically, for any point $\mathbf{x}$, we have $\mathbf{p}(\mathbf{x})=(\boldsymbol{\alpha}(\mathbf{x}),r(\mathbf{x}),m(\mathbf{x}))=f_{mat}(\mathbf{x})$, where $\boldsymbol{\alpha}\in[0,1]^{3}$ is the albedo color, $r\in[0,1]$ is the roughness, and $m\in[0,1]$ is the metallic value. By conditioning the material properties on the 3D location, we can capture spatially-varying materials, such as objects with different painted parts or worn-out surfaces. This explicit decomposition is crucial for downstream applications like material editing, a feature lacking in methods that only model view-dependent appearance (poole2022dreamfusion; lin2023magic3d).
@@ 3.2.3. Environment Lighting
We model the scene’s illumination as a distant environment map. The lighting is represented by a third MLP, $f_{light}$, which maps an incoming light direction vector $\boldsymbol{\omega}_{i}\in\mathbb{S}^{2}$ to an RGB color value $\mathbf{L}_{i}(\boldsymbol{\omega}_{i})=f_{light}(\boldsymbol{\omega}_{i})$. This formulation is flexible enough to represent complex, high-frequency lighting environments, which is a notable improvement over simpler models like spherical harmonics used in some earlier works (zhou2021nerfactor). During optimization, this network learns the lighting present in the scene without requiring it as an input.
@@ 3.3. Differentiable Rendering for Decomposition
The core of our method is a differentiable rendering process that connects our 3D scene representation to the 2D input images. We adopt the volume rendering principle from NeuS (wang2021neus) to robustly find the surface intersection and then apply a physically-based rendering equation at that surface point.
@@ 3.3.1. Ray Marching and Surface Finding
(1) $\mathbf{C}(\mathbf{r})=\int_{t_{n}}^{t_{f}}T(t)w(t)\mathbf{c}(\mathbf{r}(t),\mathbf{d})dt,$
where $t_{n},t_{f}$ are near/far bounds, $T(t)=\exp(-\int_{t_{n}}^{t}\sigma(s)ds)$ is the transmittance, $w(t)$ is a weighting function, and $\mathbf{c}(\cdot)$ is the color at point $\mathbf{r}(t)$.
Instead of a generic density $\sigma$, we use a weighting function $w(t)$ that is guaranteed to be a single, unbiased peak at the surface location. Following NeuS (wang2021neus), we define a density-like function $\sigma(t)$ from the SDF value and then derive the weight $w(t)=T(t)\sigma(t)$. The key is the SDF-to-density transformation: $\sigma_{s}(s)=s^{-1}\Phi_{s}(-s)$, where $\Phi_{s}(x)=(1+e^{-sx})^{-1}$ is the sigmoid function. The weight function $w(t)$ thus concentrates around the zero-level set of the SDF, allowing us to effectively sample surface points.
@@ 3.3.2. Physically-Based Shading
(2) $\mathbf{L}_{o}(\mathbf{x}_{s},\boldsymbol{\omega}_{o})=\int_{\Omega}f_{r}(\mathbf{x}_{s},\boldsymbol{\omega}_{i},\boldsymbol{\omega}_{o})\mathbf{L}_{i}(\boldsymbol{\omega}_{i})(\mathbf{n}(\mathbf{x}_{s})\cdot\boldsymbol{\omega}_{i})d\boldsymbol{\omega}_{i},$
@@
The terms $D,F,G$ are standard functions of the surface roughness $r$, view/light directions, and the Fresnel term $F$ additionally depends on the metallic property $m$ and albedo $\boldsymbol{\alpha}$ (munkberg2022nvdiffrec). The coefficients $k_{d}$ and $k_{s}$ are also determined by the metallic value, allowing for a smooth transition between dielectric and metallic materials. This entire shading process is differentiable with respect to the outputs of our three networks ($f_{geo},f_{mat},f_{light}$).
@@ 3.4. Optimization and Loss Functions
The networks are trained end-to-end by minimizing a composite loss function. The ability to learn from complex, multi-modal, and sometimes noisy data is crucial, a challenge shared by fields as diverse as sign language production and WiFi-based sensing (wang2024targetoriented).
@@ 3.4.1. Photometric Reconstruction Loss
The primary supervision signal is the photometric reconstruction loss, which measures the difference between the rendered pixel colors $\mathbf{C}(\mathbf{r})$ and the ground-truth colors from the input images $\mathbf{C}_{gt}(\mathbf{r})$. We use a combination of an L1 loss and the LPIPS perceptual loss (srinivasan2023nero):
[…]
@@ 3.4.2. Geometric Regularization
[…]
@@
We also use a mask loss to enforce that SDF values inside the object (as defined by the input masks) are negative. This robust supervision is critical, much like how robust authentication is needed to prevent attacks in other sensing domains (huang2024phyfinatt).
@@ 3.4.3. Material and Lighting Priors
[…]
@@
Additionally, we add a sparsity prior on the metallic parameter to encourage it to be close to either 0 (dielectric) or 1 (metal), reflecting the behavior of most common materials. We also regularize the total intensity and smoothness of the predicted environment light map from $f_{light}$ to prevent overly complex lighting from being ”baked” into the material albedo. The design of these priors is crucial for stable training, a principle that holds true when designing complex hierarchical systems, whether for federated learning (liu2024finch) or visual understanding .
@@ 3.5. Implementation Details
Our geometry and material networks are implemented as MLPs with 8 layers, 256 hidden units, and sinusoidal positional encodings (mildenhall2020nerf) for the input coordinates to capture high-frequency details. The lighting network is a smaller 4-layer MLP. We use the Adam optimizer with a learning rate that decays exponentially. The entire framework is implemented in PyTorch. The training process, summarized in Algorithm 1, typically converges after 200k iterations on a single NVIDIA A100 GPU. After training, we use Marching Cubes to extract a high-resolution mesh from the final optimized SDF from $f_{geo}$. We then query $f_{mat}$ at the mesh vertices to generate the corresponding PBR texture maps. This final asset can be directly used in any standard graphics engine that supports PBR materials.
@@ 4.1.1. Datasets
We evaluate our method on both synthetic and real-world datasets to demonstrate its effectiveness and generalization capabilities. The diversity of data is crucial for robust evaluation, a principle that is vital across various domains, from training models for sign language to ensuring robust performance in WiFi-based activity recognition under interference (huang2023phaseanti).
@@
For quantitative evaluation of both geometry and material decomposition, we use a synthetic dataset inspired by the one used in NVDIFFREC (munkberg2022nvdiffrec). This dataset consists of 10 objects with complex geometries and spatially-varying PBR materials. Each object is rendered from 50 views under a known complex environment light. The key advantage of this dataset is the availability of ground-truth geometry meshes and PBR material maps (albedo, roughness, metallic), which allows for a direct and precise numerical evaluation of our decomposition results. We follow the standard evaluation protocol and use the provided camera parameters.
@@
To validate our method on real-world data, we use the widely-adopted DTU dataset . The DTU dataset contains multi-view images of various objects captured under controlled laboratory conditions with a fixed camera rig and several different lighting conditions. This dataset is a standard benchmark for multi-view surface reconstruction (wang2021neus; yariv2021volsdf). Since ground-truth material properties are not available for these real objects, evaluation is primarily focused on the quality of the geometric reconstruction and the visual fidelity of novel view synthesis and relighting results. The challenge of working with real-world data, where ground truth is scarce, mirrors challenges in other sensing domains that rely on unsupervised or weakly-supervised methods, such as gesture recognition from commodity signals (gu2023attentionbased).
@@ 4.1.2. Evaluation Metrics
We adopt a comprehensive set of metrics to evaluate different aspects of our framework’s performance. The choice of metrics is critical for a fair comparison, reflecting the multi-faceted nature of the inverse rendering problem, much like how multi-objective optimization is key for efficient model compression .
@@
  • Novel View Synthesis: For both datasets, we evaluate the quality of rendered novel views against ground-truth images using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and the Learned Perceptual Image Patch Similarity (LPIPS) (srinivasan2023nero). PSNR and SSIM measure pixel-level fidelity, while LPIPS better correlates with human perceptual judgment, a crucial aspect for tasks aiming at human-centric understanding, such as visual dialog or emotion recognition (gu2023wife).
@@ 4.1.3. Baselines and Implementation Details
  • NeuS (wang2021neus): A state-of-the-art method for high-quality neural surface reconstruction. This baseline represents methods that focus solely on geometry.
  • NVDIFFREC (munkberg2022nvdiffrec): A leading method for inverse rendering that directly optimizes a mesh representation. It is a strong baseline for joint geometry and material optimization.
  • Mat-NeuS (zhang2022matneus): An extension of NeuS that incorporates material decomposition, making it a very relevant baseline for our approach.
  • IRON (zhang2022iron): Another powerful inverse rendering framework that demonstrates high-quality object reconstruction and material estimation.
@@
Our framework is implemented in PyTorch. All experiments are conducted on NVIDIA A100 GPUs with 40GB of memory. We use the Adam optimizer with an initial learning rate of $5\times 10^{-4}$, which decays exponentially over the course of training. Each model is trained for 300k iterations with a batch size of 512 rays. The loss weights are set empirically, with $\lambda_{L1}=1.0$, $\lambda_{LPIPS}=0.2$, $\lambda_{eikonal}=0.1$, and $\lambda_{mat}=0.01$. The design of an efficient and stable training pipeline is paramount, a concern shared by those developing complex distributed systems for tasks like federated learning (liu2024finch; liu2024federated) or dealing with the security implications of advanced sensing technologies (huang2024keystrokesniffer). Our code and pre-trained models will be made publicly available to facilitate future research.
@@ 4.2. Quantitative Comparison
Table 1 provides a detailed breakdown of performance on the synthetic dataset, where ground-truth geometry and materials are available for precise evaluation. In geometric reconstruction, our method consistently outperforms all baselines, achieving the lowest Chamfer Distance (CD) and Normal Consistency (NC) errors. Compared to NeuS (wang2021neus), which focuses solely on geometry, our joint optimization of shape and material yields a more accurate surface. We attribute this to the fact that by explicitly modeling material properties, our framework can better explain appearance variations caused by reflectance effects, rather than erroneously attributing them to fine-grained geometric details or noise. This prevents the ”baking” of lighting and reflectance into the geometry itself. Our geometric fidelity even surpasses that of leading inverse rendering methods like NVDIFFREC (munkberg2022nvdiffrec) and Mat-NeuS (zhang2022matneus) which we believe is due to our robust lighting model and the comprehensive set of physical and geometric priors that regularize the optimization process. The challenge of disentangling correlated factors is a common theme in machine perception, where, for instance, separating an event signal from noise is key in audio-visual analysis .
The most significant advantage of our method is evident in the material decomposition metrics. As shown in Table 1, MatDecompSDF achieves a substantially higher PSNR across all PBR maps (albedo, roughness, and metallic) compared to all baselines. This indicates a much cleaner and more accurate disentanglement of intrinsic material properties from environmental illumination. While methods like IRON (zhang2022iron) can recover plausible materials, they often exhibit minor artifacts, such as residual lighting baked into the albedo map. Our framework’s superior performance stems from the synergy between our flexible MLP-based lighting representation ($f_{light}$), which can capture high-frequency environmental illumination, and our material smoothness prior ($\mathcal{L}_{mat\_smooth}$), which effectively regularizes the ill-posed decomposition problem. This demonstrates that a well-constrained optimization, guided by physical principles, is paramount, a lesson that holds true in fields like model compression where multi-objective approaches are needed to balance competing goals .
@@
Table 1. Quantitative comparison on the synthetic dataset. We report Chamfer Distance (CD, $m$, $\downarrow$), Normal Consistency (NC, degrees, $\downarrow$), and PSNR (dB, $\uparrow$) for material maps and rendering quality. Best results are in bold.
@@
NeuS (wang2021neus) 0.85 16.2 - - - 29.81 -
NVDIFFREC (munkberg2022nvdiffrec) 0.79 15.8 25.12 23.45 24.88 31.23 28.95
Mat-NeuS (zhang2022matneus) 0.75 15.1 26.03 24.11 25.01 31.55 29.24
IRON (zhang2022iron) 0.81 15.5 25.88 23.97 24.92 31.40 29.08
@@
Table 2 shows the performance on the challenging real-world DTU dataset. Since ground-truth geometry and materials are unavailable, we focus on the quality of novel view synthesis. Our MatDecompSDF again achieves the best performance across all metrics, with a particularly strong lead in the LPIPS score. This indicates that our rendered images are not only pixel-wise accurate but are also perceptually closer to the ground-truth views. We attribute this to two factors: first, the inclusion of the $\mathcal{L}_{LPIPS}$ term in our photometric loss guides the optimization towards perceptually plausible solutions, a strategy whose importance is recognized in generative modeling (srinivasan2023nero). Second, by modeling the scene with physically accurate materials and lighting, our framework produces more realistic shading and highlights, which are key components of perceptual quality. This is in contrast to geometry-only methods like NeuS, whose view-dependent appearance model can sometimes produce results that look flat or lack realism under complex lighting. The ability to perform robustly on real-world data with its inherent noise and complexities is a testament to our framework’s design, echoing the need for robust systems in other real-world sensing applications, from federated learning in heterogeneous networks (liu2024federated) to secure WiFi-based gesture recognition (gu2023attentionbased).
@@
NeuS (wang2021neus) 25.31 0.871 0.145
NVDIFFREC (munkberg2022nvdiffrec) 25.88 0.885 0.131
Mat-NeuS (zhang2022matneus) 26.05 0.892 0.124
IRON (zhang2022iron) 25.97 0.888 0.129